OpenStudy (anonymous):

sqrt 5x^3 x sqrt 5x^5

OpenStudy (anonymous):

\[\sqrt{25x^8}\] = \[5x^{4}\]

OpenStudy (anonymous):

\[\sqrt{5x^3}*\sqrt{5x^5}=\sqrt{5}*\sqrt{x^3}*\sqrt{5}*\sqrt{x^5}=5*\sqrt{x^8}=5x^4\]